Find the coordinates of the point on the curve y=2x+3 that is closest to the point (3.0).If f(0) = sine, find f (2/6).
Answer #1

We'll answer the first question since the exact one wasn't specified. Please submit a new question specifying the one you'd like answered.

* Cx) = 818-3) + 2(2x+3). 2 4s 2x+3 and point (3,0) 22+ 3 het (any) be the point on the curve. Now distance between carg) andY= 2 x+ 3 y = 2(-0.6) + 3 = -1:2+3 = 1.8 Hence closest point us C -0.6,1.8)